# Look Twice Before You Answer: Memory-Space Visual Retracing for Hallucination Mitigation in Multimodal Large Language Models

### v1

@@ 3 Methodology
In this section, we first formulate the generation process of MLLMs to facilitate a clearer understanding of our MemVR. Moreover, we introduce our hypothesis about the causes of hallucinations and discuss visual retracing and its dynamic strategy. Further, we conduct theoretical analysis.
3.1 MLLMs Generation Process
Typically, MLLMs are composed of a visual encoder ${C}_{v}$, a text embedding layer, $L$ number of transformer layers, and an affine layer $\zeta(\cdot)$ which predicts the distribution of the next token, with diverse modalities as inputs, e.g., images and text. Regardless of specific architectural variations, MLLMs commonly employ ${C}_{v}$ to extract vision tokens from the raw image and project it into modality-shared feature space via an MLP or Q-Former module (Wadekar et al., 2024). Aligned vision tokens are represented as $X_{v}=\{x_{1},x_{2},\dots,x_{n_{v}}\}$ and used as part of the LLM input alongside the text tokens $X_{q}=\{x_{n_{v}+1},x_{n_{v}+2},\dots,x_{n_{v}+n_{q}-1}\}$ that are embedded from tokenized text input by the embedding layer. Subsequently, the vision and text tokens are concatenated as the final input sequence and we denote it as $\{x_{i}\}_{1}^{t_{n}-1}$ where $t_{n}=n_{v}+n_{q}$, which is then fed into successive transformer layers. We denote the output of the $l$-th layer as $h_{t}^{(l)}$. Then, the vocabulary head $\zeta(\cdot)$ is used to predict the probability of the next token $x_{t}$ among the vocabulary set $\mathcal{X}$ as follows,
$p(x_{t}\mid x_{<t})=\mathrm{softmax}(\zeta(h_{t}^{(L)}))_{x_{t}},x_{t}\in\mathcal{X}.$ (1)
3.2 What Causes Hallucinations
A hypothesis on the cause of hallucinations. Informed by the phenomenon of catastrophic forgetting (Zhai et al., 2023) in MLLMs, we argue that the capabilities of LLMs to comprehend and memorize different modalities are quite distinct. Taking image and text inputs as an example, since an image possesses a much higher information density than a piece of text, it is reasonable to assume that LLMs struggle to understand and memorize vision tokens compared to text tokens, prone to fantasies.
Uncertainty quantification. Following the DoLa (Chuang et al., 2023), we compute the probability of the next token via the vocabulary head $\zeta$ on each layer during reasoning. Then, we introduce an entropy-based metric (Farquhar et al., 2024) to quantify the output uncertainty as $u=\sum-p_{k}\log p_{k}/\log{K}$, where $\{p_{k}\}_{i=1}^{K}$. With uncertainty, we make an important assumption.
Assumption 3.1.
LUFH: The Lower the Uncertainty, the Fewer the Hallucinations to be generated.
We define $\gamma$ as the threshold that separates high uncertainty from low uncertainty in predictions. The candidate layer with uncertainty exceeding $\gamma$ is termed a premature layer. The proof of Assumption 3.1 (LUFH) is provided in Section 4. Based on the hypothesis, we aim to complement visual evidence in MLLMs to eliminate the hallucination caused by visual forgetting. In the following, we discuss our motivation, how MemVR is implemented, and why it can work in Section 3.3 and 3.4.
3.3 Relationship Between Hallucinations and Uncertainty
As findings of Chen et al. (2024) in LLMs: “incorrect tokens generally exhibit higher entropy than correct ones”, we also observe this phenomenon in MLLMs (visualization cases are shown in Appendix C.3). This implies the effectiveness of entropy-based metrics for detecting hallucinations. In this work, we use uncertainty as the metric. We present our in-depth findings in this section.
Figure 3: Uncertainty of different early layers to predict the next token. Rows denote indices of the early layers, and column names are decoded tokens in each step. Uncertainty distribution is dynamic.
Finding #1: In the context of tokens involving objects, attributes or relations, uncertainty is high.
We conduct preliminary analysis with 32-layer LLaVA-1.5-7B. Specifically, we compute the uncertainty in the output distributions $p^{(l)}(\cdot\mid x_{<t})$ of early exiting layers. Figure 3 shows the uncertainty scores of different early layers when decoding the answer, we can observe that the computed uncertainty remains relatively high in later layers when predicting key entity objects, attributes, or relations, such as wooden, filled, and food in Figure 3. This phenomenon suggests that LLM is still uncertain about its predictions in the last few layers and may inject more factual knowledge into the predictions. On the other hand, when predicting function words and those tokens copied from the question, e.g., image, a, with, we observe that the uncertainty becomes very low from the middle layers. This finding implies that the model is deterministic for easy-to-predict tokens at the intermediate layer and keeps the distribution of outputs almost constant at higher layers, however, it is more uncertain for difficult-to-predict key tokens and may constantly change its predictions until the final layer.
Finding #2: The uncertainty distribution of different tokens is dynamic during the generation.
As can be seen in Figure 3, when visual factual knowledge is required for the prediction of the next token, such as wodden, filled, food, most layers of LLaVA are of higher uncertainty, and appear to shift the prediction of the later layers. The prediction of the next tokens such as quantity, target category, color, relation, etc. requires evident visual knowledge, but not for function words or commas
From a qualitative perspective, intermediate values of uncertainty reflect the thinking or revision process of MLLMs, while too high uncertainty implies that the model is confused and relies on random guesses. For complex input images, the uncertainty of the output distributions tends to rise, which can propagate through subsequent layers, resulting in incorrect predictions, or gibberish.
Figure 4: The illustration of how dynamic premature layer injection works.
3.4 Visual Retracing and Its Dynamics
Motivated by the above findings, we propose re-injecting visual evidence during elevated uncertainty in the model’s reasoning. This strategy treats visual tokens as anchors to recalibrate off-target predictions and reduces uncertainties in object, attribute, relationship tokens. Experimental results also demonstrate that our method reduces uncertainty and alleviates hallucinations as shown in Figure 9. We term this schema of re-injecting visual evidence as “visual retracing” that is to be elaborated further in Section 3.4.1. Further, we design a dynamic injection strategy, detailed in Section 3.4.2, ensuring timely visual evidence when generating uncertain visual-reliant tokens (Figure 4).
3.4.1 FFN with Visual Retracing
We introduce the implementation of our proposed memory-space visual retracing method. Previous study (Geva et al., 2021) has found that FFN acts as a key-value memory storing factual knowledge. Inspired by the fact that FFN executes analogous retrieval from its key-value memory, we consider “visual retracing” to serve as a simplified and efficient information re-retrieval process (Jie et al., 2024). Concretely, given an hidden token $x\in\mathbb{R}^{d}$ and vision features $z_{v}=(z_{v,1},\dots,z_{v,n_{v}})^{\top}\in\mathbb{R}^{n_{v}\times d^{\prime}}$, $\mathcal{K}_{\propto}\left({z}_{v,i}\right),\mathcal{V}_{\propto}\left({z}_{v,i}\right)\in\mathbb{R}^{d}$ are the key and value corresponding to ${z}_{v,i}$, visual retracing follows,
${\operatorname{FFN}}_{\circlearrowright}^{(l)}({x}\propto z_{v})=(1-\alpha)\operatorname{FFN}^{(l)}(x)+\alpha\underline{\Lambda}_{\circlearrowright}^{(l)}({z}_{v}\mid{x}),$ (2)
where $\alpha$ denotes the injection ratio of visual memory (proportional to image complexity), FFN comprises two FC layers with non-linear activation in between and can be formulated as $\operatorname{FFN}({x})=\phi\left({x}{W}_{1}\right){W}_{2}^{\top}$, $\phi$ is activation function like ReLU or SiLU (Liu et al., 2020). Separately, the weight matrices can be rewritten as: ${W}_{1}=\left({k}_{1},{k}_{2},\ldots,{k}_{m}\right),{W}_{2}=\left({v}_{1},{v}_{2},\ldots,{v}_{m}\right)^{\top}$, in which ${k}_{i}\in\mathbb{R}^{d}$ and ${v}_{i}\in\mathbb{R}^{d}$ are entries of key and value, respectively. Thus, FFN can be interpreted as using input ${x}$ as the query to calculate its similarity with keys to search for matching values. Analogously, we consider a simple and efficient retrieval process for visual retracing on $l$-th premature layer as,
@@ 3.4.1 FFN with Visual Retracing
$\underline{\Lambda}_{\circlearrowright}^{(l)}({z}_{v}\mid{x})=\sum\nolimits_{i=1}^{n_{v}}\phi\left(\left\langle{x},\mathcal{K}_{\propto}\left({z}_{v,i}\right)\right\rangle\right)\cdot\mathcal{V}_{\propto}\left({z}_{v,i}\right).$ (3)
From the perspective of FFN, we regard the $\left(\mathcal{K}_{\propto}\left({z}_{v,i}\right),\mathcal{V}_{\propto}\left({z}_{v,i}\right)\right)$ as new key-value entries to supplement vision-related evidence. $\mathcal{K}_{\propto}$ and $\mathcal{V}_{\propto}$ are key functions for aligning the dimension between vision features $z_{v}$ and the hidden token $x$. In MLLMs, $\mathcal{K}_{\propto}$ and $\mathcal{V}_{\propto}$ are regarded as visual projectors, e.g., MLPs or Q-Former (Li et al., 2023a), for aligning features of different modalities.
3.4.2 Dynamic Premature Layer Injection
@@
To magnify the effectiveness of visual retracing, the optimal premature layer should ideally be the layer most uncertain about probable answers to visual questions. In practice, we consider that the uncertainty of a candidate layer exceeding the threshold $\gamma$ warrants visual retracing. Inspired by the fact that early exit patterns (Teerapittayanon et al., 2016; Elbayad et al., 2020; Schuster et al., 2022) have proven effective in directly employing the language heads $\zeta$ to the hidden states of the middle layers, even without a special training process (Kao et al., 2020), we compute the uncertainty of the next token probability on the early layers for reasoning. As our Finding #2, we utilize layer-specific uncertainty to allow for dynamic premature layer injection at each time step as illustrated in Figure 3.
Algorithm 1 Dynamic Injection Strategy
1:$h_{t},{z}_{v}$ denote hidden states and visual evidence. We set trigger = True.
2:for $l$ = 1 to $L-1$ do
3:     $p^{(l)}=\mathrm{softmax}(\zeta(h_{t}^{(l)}))_{x_{t}}$
4:     $u^{(l)}={\sum-p^{(l)}\log p^{(l)}}/{\log{K}}.$
@@
6:         Execute $\underline{\Lambda}_{\circlearrowright}^{(l+1)}({z}_{v}\mid{h_{t}^{(l+1)}})$
7:         Select ${\operatorname{FFN}}_{\circlearrowright}^{(l+1)}({h_{t}^{(l+1)}}\propto z_{v})$
8:         trigger=False # only once
9:     else
10:         Select ${\operatorname{FFN}}^{(l+1)}({h_{t}^{(l+1)}})$
11:     end if
12:end for
For MLLMs with different numbers of layers, we partition the layers into several (typically three) buckets according to the total layers for sensible visual retracing triggers, as detailed in Appendix C.1. This dynamic injection strategy identifies desirable premature layers for visual retracing based on output uncertainty, thus better leveraging insights from different layers.
Static Fixed Layer Injection. In addition to the dynamic premature layer injection strategy, another more straightforward strategy worth considering is to perform a brute-force experiment on all possible early layers using a validation set and selecting the layer with the best average performance. We refer to this simple strategy as MemVR-static. However, the MemVR-static approach presents two limitations: (I) it requires more extensive hyperparameter tuning across layers, and (II) the optimal layer is highly sensitive to data distribution, necessitating an in-distribution validation set. In contrast, our proposed dynamic layer injection strategy mitigates these challenges by reducing the layer search space and improving robustness without depending on in-distribution validation. Empirical comparisons between our MemVR using dynamic and static strategies are provided in Section 4.2 and Table 4.
3.5 Theoretically Understanding Why MemVR Works
In order to gain further insight into the reasons behind the effectiveness of MemVR in mitigating hallucinations and its robust performance on general benchmarks, we attempt to explain these phenomena via the following three theorems from an information-theoretic perspective.
@@ 3.5 Theoretically Understanding Why MemVR Works
The reinjection of $Z_{v}$ at intermediate layers of the model facilitates the replenishment of critical visual information, which may have been lost or distorted through earlier layers. This process increases MI between the hidden states and the visual features, ensuring adequate visual context is preserved.
@@
Let $Y$ be the target output dependent on hidden states. If MI between ${H}_{vq}$ and ${Z}_{v}$ increases, then conditional entropy $H(Y\mid{H}^{(l)}_{vq})$ decreases with $H(Y\mid\hat{H}_{vq})\leq H(Y\mid H_{vq})$.
According to Theorem 3.1, 3.2, and DPI (Cover et al., 1991), MemVR improves the quality of the representations $H_{vq}$ and reduces the uncertainty in the output $Y$. As a result, the probability of hallucinations decreases, which is consistent with the observed findings of hallucination mitigation.
@@
Within the Information Bottleneck (IB) framework, the loss of objective function, represented by the notation $\mathcal{L}(T)$, is optimized by MemVR, which is defined as $\mathcal{L}(\hat{H}_{vq})\leq\mathcal{L}(H_{vq})$, where $\mathcal{L}(H_{vq})=I(H_{vq};X_{vq})-\beta I(H_{vq};Y)$ is IB loss, and $\ beta$ is a trade-off parameter.
Anchored in the information bottleneck framework, MemVR optimizes the delicate balance between retaining relevant information from multimodal inputs and compressing non-essential details, thereby safeguarding the predictive performance of the hidden representations for the target output $Y$.
The theoretical underpinnings of MemVR are supported by the Data Processing Inequality (Cover et al., 1991) and the contraction properties of stochastic mappings in deep neural networks, as shown in various studies on the Information Bottleneck Principle (Achille & Soatto, 2018). By enhancing mutual information and reducing the uncertainty in hidden states, MemVR effectively mitigates hallucinations while preserving computational efficiency. Detailed proofs are in Appendix B.
Method   MSCOCO   A-OKVQA   GQA
%Accuracy %F1 Score %Accuracy %F1 Score %Accuracy %F1 Score
LLaVA1.5-7B 81.38  $\uparrow$0.0 79.65  $\uparrow$0.0 79.13  $\uparrow$0.0 79.10  $\uparrow$0.0 79.00  $\uparrow$0.0 79.13  $\uparrow$0.0
+ VCD (Leng et al., 2024) 84.66  $\uparrow$3.3 84.52  $\uparrow$4.9 80.99  $\uparrow$1.8 82.30  $\uparrow$3.2 81.74  $\uparrow$2.7 82.16  $\uparrow$3.0
+ OPERA (Huang et al., 2024a) 84.77  $\uparrow$3.4 85.46  $\uparrow$5.8 84.27  $\uparrow$5.1 84.08  $\uparrow$5.0 84.03  $\uparrow$5.0 83.83  $\uparrow$4.7
+ MemVR (Ours) 87.00  $\uparrow$5.7 85.87  $\uparrow$6.2 86.21  $\uparrow$7.0 86.64  $\uparrow$7.5 85.25  $\uparrow$6.2 85.59  $\uparrow$6.4
Qwen-VL-10B 83.79  $\uparrow$0.0 81.13  $\uparrow$0.0 84.74  $\uparrow$0.0 83.27  $\uparrow$0.0 84.41  $\uparrow$0.0 82.66  $\uparrow$0.0
+ VCD (Leng et al., 2024) 84.27  $\uparrow$0.4 82.12  $\uparrow$1.0 84.09  $\downarrow$0.7 82.53  $\downarrow$0.7 83.73  $\downarrow$0.7 82.75  $\uparrow$0.1
+ OPERA (Huang et al., 2024a) 84.93  $\uparrow$1.1 83.41  $\uparrow$2.3 - - - -
+ MemVR (Ours) 84.07  $\uparrow$0.3 81.55  $\uparrow$0.4 86.43  $\uparrow$1.8 85.56  $\uparrow$2.3 85.69  $\uparrow$1.3 84.53  $\uparrow$1.9
Table 2: Performance evaluation on POPE. The best results in each scenario are bolded for clarity. We report the averages under the three settings, e.g., Random, Popular, and Adversarial to show the robustness of the different methods directly. Green denotes improvement, and Red means degradation.
Method Commonsense QA(Reasoning)   Object-level Hallucination   Attribute-level Hallucination  Total Scores
Existence Count Position Color
LLaVA1.5-7B 110.71  $\uparrow$0.0 175.67  $\uparrow$0.0 124.67  $\uparrow$0.0 114.00  $\uparrow$0.0 151.00  $\uparrow$0.0 676.05
+ VCD (Leng et al., 2024) 112.86  $\uparrow$9.9 184.66  $\uparrow$9.0 138.33  $\uparrow$13.6 128.67  $\uparrow$14.6 153.00  $\uparrow$2.0 717.52
+ OPERA (Huang et al., 2024a) 115.71  $\uparrow$5.5 180.67  $\uparrow$5.0 133.33  $\uparrow$8.6 123.33  $\uparrow$9.3 155.00  $\uparrow$4.0 708.04
+ MemVR (Ours) 121.42  $\uparrow$18.5 190.00  $\uparrow$14.3 155.00  $\uparrow$30.3 133.33  $\uparrow$19.3 170.00  $\uparrow$19.0 769.75
Qwen-VL-10B 106.40  $\uparrow$0.0 155.00  $\uparrow$0.0 127.67  $\uparrow$0.0 131.67  $\uparrow$0.0 173.00  $\uparrow$0.0 693.74
+ VCD (Leng et al., 2024) 104.33  $\downarrow$2.1 156.00  $\uparrow$1.0 131.00  $\uparrow$3.3 128.00  $\downarrow$3.6 181.67  $\uparrow$8.6 701.00
+ OPERA (Huang et al., 2024a) 104.33  $\uparrow$2.2 165.00  $\uparrow$6.9 145.00  $\uparrow$4.8 133.33  $\uparrow$1.6 180.00  $\uparrow$7.0 727.66
+ MemVR (Ours) 120.00  $\uparrow$13.6 185.00  $\uparrow$30.0 145.00  $\uparrow$17.3 123.33  $\downarrow$8.3 185.00  $\uparrow$12.0 758.33
Table 3: Results on the hallucination subset of MME (including commonsense reasoning, existence, count, position, color scores). The best are in bold. MemVR achieves dramatic improvements.

### v2

@@ 4 Methodology
Inspired by the common cognition: when the memory of an image seen the moment before is forgotten, people will look at it again for factual answers, we design look-twice mechanism, which treats visual tokens as supplementary evidences, re-injecting them into MLLMs through FFN as “key-value memory” at the middle trigger layer. This look-twice mechanism occurs when the model exhibits high uncertainty during inference, effectively enhancing factual alignment and modality balance. All details are listed below.
4.1 Preliminary: Reformulation of FFN
Vanilla FFN comprises two fully connected layers with non-linear activation in between. We suppose $\boldsymbol{x}\in\mathbb{R}^{d}$ as an input token of the FFN, and FFN function can be formulated as
$\operatorname{FFN}(\boldsymbol{x})=\phi\left(\boldsymbol{x}\boldsymbol{W}_{1}\right)\boldsymbol{W}_{2}^{\top},$ (2)
where $\phi$ is activation function like ReLU or SiLU (Liu et al., 2020), and $\boldsymbol{W}_{1},\boldsymbol{W}_{2}\in\mathbb{R}^{d\times D}$ are the weight matrices, in usual $D=4d$. Peculiarly, $\boldsymbol{W}_{1}$ and $\boldsymbol{W}_{2}$ can be rewritten as
$\boldsymbol{W}_{1}=(\boldsymbol{k}_{1},\boldsymbol{k}_{2},\ldots,\boldsymbol{k}_{D}),\boldsymbol{W}_{2}=(\boldsymbol{v}_{1},\boldsymbol{v}_{2},\ldots,\boldsymbol{v}_{D}),$ (3)
where $\boldsymbol{k}_{i},\boldsymbol{v}_{i}\in\mathbb{R}^{d}$ denote entries of key and value, respectively. As a result, the FFN can be reformulated as
$\operatorname{FFN}(\boldsymbol{x})=\sum\phi\left(\left\langle\boldsymbol{x},\boldsymbol{k}_{i}\right\rangle\right)\cdot\boldsymbol{v}_{i}\;.$ (4)
Thus, the FFN function can be construed as using input $\boldsymbol{x}$ as a query to measure similarity with keys, find matching values, and gather values by similarity, which works like a key-value memory storing the factual knowledge as found in previous studies (Geva et al., 2021; Jie et al., 2024).
4.2 FFN with Visual Retracing
Motivated by the findings in section §3.2 and §3.4, we propose Visual Retracing (VR), i.e., reinjecting visual evidence into the middle layer of the text decoder during elevated uncertainty during reasoning. This strategy treats visual tokens as anchors to recalibrate off-target predictions and reduces uncertainties in object, attribute, relationship tokens. Experimental results also demonstrate that our method reduces uncertainty and alleviates hallucinations as shown in Figure 8. The reason we call this pattern of reinjecting visual evidence “visual retracing” is that the model finds and refreshes key visual memories based on the hidden states in this process. In particular, inspired by the fact that FFN executes analogous retrieval from its key-value memory, we consider VR to serve as a simplified and efficient information re-retrieval process. Given a hidden token $\boldsymbol{x}\in\mathbb{R}^{d}$ and dimension-aligned vision tokens $\boldsymbol{z}_{v}$, FFN with visual retracing at $l$-th layer can be written as follows
${\operatorname{FFN}}^{(l)}(\boldsymbol{x}\propto\boldsymbol{z}_{v})=\alpha\underline{\Delta}+(1-\alpha)\operatorname{FFN}^{(l)}(\boldsymbol{x}),$ (5)
where $\boldsymbol{z}_{v}=(\boldsymbol{z}_{v,1},\dots,\boldsymbol{z}_{v,N_{v}})\in\mathbb{R}^{d\times N_{v}}$, ${x}\propto\boldsymbol{z}_{v}$ denotes execute VR $\underline{\Delta}$ from $\boldsymbol{x}$ to visual features $\boldsymbol{z}_{v}$, and $\alpha\in[0,1]$ denotes injection ratio of visual memory through the FFN layer which proportional to image complexity. Specifically, instead of performing retrieval via cross-attention layers as in previous approaches (Li et al., 2022; Alayrac et al., 2022), we consider a simple retrieval process for VR as,
@@ 4.2 FFN with Visual Retracing
$\underline{\Delta}({\boldsymbol{z}}_{v}\mid\boldsymbol{x})=\sum\nolimits_{i=1}^{N_{v}}\phi(\left\langle\boldsymbol{x},{\boldsymbol{z}}_{v,i}\right\rangle)\cdot{\boldsymbol{z}}_{v,i}.$ (6)
From the perspective of FFN, visual retracing works by treating $\boldsymbol{x}$ as a query, and $\left\langle{\boldsymbol{z}}_{v,i}:{\boldsymbol{z}}_{v,i}\right\rangle$ as new key-value entries (visual evidence) to supplement vision-related information in the hidden states. In this information re-retrieval process, MemVR does not introduce any parameters that need to be trained. Notably, since the size of key-value memory $D$ in FFN typically far exceeds the number of visual tokens $N_{v}$ (for instance, $D=11008$ in LLaMA-7B and $N_{v}=256$ for ViT-L/14, $N_{v}\ll D$), the computation of MemVR is negligible. Thus, VR operation is more efficient than the cross-attention mechanism with quadratic complexity.
4.3 Dynamic Triggered Visual Retracing
@@
To magnify the effectiveness of VR, the optimal strategy should be to trigger VR dynamically based on the token, and the selection of the trigger layer should also be dynamically determined. In practice, we consider that the uncertainty of a candidate layer exceeding the threshold $\gamma$ warrants visual retracing. Inspired by the fact that early exit patterns (Elbayad et al., 2020; Schuster et al., 2022) have proven effective in directly employing the language heads $\zeta$ to the hidden states of the middle layers, even without a special training process (Kao et al., 2020), we compute the uncertainty of the next token probability on the early layers for reasoning. We utilize layer-specific uncertainty to allow for dynamic premature layer injection at each time step.
Dynamic Triggered MemVR. For MLLMs with different numbers of layers, as Algorithm 1 shown, the dynamic triggered MemVR strategy called MemVR-dynamic, identifies desirable premature layers among the candidate layers for visual retracing based on output uncertainty of different layers, thus better amplifying the effect of visual retracing.
Algorithm 1 Dynamic Triggered MemVR Strategy
0:  MLLM $\mathcal{M}_{\theta}^{\text{MLLM}}$, text query $x$, image input $v$.
0:  Model response $y_{t}^{b}$.
1:  At every decoding step $t$:
2:  Initial set to trigger = True.
3:  for $l$ = 1 to $L-1$ do
4:     $u^{(l)}={\sum-p_{\theta}^{(l)}\log p_{\theta}^{(l)}}/{\log{N}}.$   % uncertainty
@@
6:        Execute $\underline{\Delta}({\boldsymbol{z}}_{v}\mid{h_{t}^{(l+1)}})$4, Eq. (6)}
7:        Select ${\operatorname{FFN}}^{(l+1)}({h_{t}^{(l+1)}}\propto\boldsymbol{z}_{v})$4, Eq. (5)}
8:        trigger = False   % set to only look-twice
9:     end if
10:  end for
11:  $\mathcal{M}_{\theta}^{\text{MLLM}}$ decoding, obtain current token $y_{t}^{b}$.
Static Triggered MemVR. Besides MemVR-dynamic, another more straightforward strategy worth considering is to perform a brute-force experiment on all possible early layers using a validation set and selecting the layer with the best average performance. We refer to this simple strategy as MemVR-static. Nonetheless, MemVR-static exhibits two notable limitations. Firstly, it demands more comprehensive hyperparameter tuning across different layers. Secondly, the optimal layer is extremely sensitive to the data distribution, which mandates the use of an in-distribution validation set.
MemVR-dynamic vs. -static. In contrast to MemVR-static, MemVR-dynamic strategy effectively alleviates these challenges. It achieves this by narrowing down the layer search space and enhancing robustness, all without relying on an in-distribution validation set. Empirical comparisons between the proposed MemVR implemented with the dynamic and static strategies are presented in Section §5.4 and Table 10.
4.4 Theoretical Analysis
To further understand why MemVR effectively mitigates hallucinations and performs robustly on general benchmarks, we explain these phenomena using three theorems below.
@@ 4.4 Theoretical Analysis
Let $\boldsymbol{x}$ be the hidden states of FFN and $\hat{\boldsymbol{x}}$ be after reinjecting visual evidence $\boldsymbol{z}_{v}$. MemVR enhances Mutual Information (MI) between $\hat{\boldsymbol{x}}$ and $\boldsymbol{z}_{v}$ as:
$I(\hat{\boldsymbol{x}};\boldsymbol{z}_{v})\geq I(\boldsymbol{x};\boldsymbol{z}_{v}).$ (7)
@@
Let $\boldsymbol{y}$ be the target output dependent on hidden states. If MI between $\boldsymbol{x}$ and $\boldsymbol{z}_{v}$ increases, then conditional entropy $H(\boldsymbol{y}\mid\boldsymbol{x})$ decreases with
$H(\boldsymbol{y}\mid\hat{\boldsymbol{x}})\leq H(\boldsymbol{y}\mid\boldsymbol{x}).$ (8)
@@
Within the Information Bottleneck (IB) framework, the loss of objective function, represented by the notation $\mathcal{L}(T)$, is optimized by MemVR, which is defined as $\mathcal{L}(\hat{\boldsymbol{x}})\leq\mathcal{L}(\boldsymbol{x})$, where $\mathcal{L}(\boldsymbol{x})=I(\boldsymbol{x};\boldsymbol{c})-\beta I(\boldsymbol{x};\boldsymbol{y})$ is IB loss, $\boldsymbol{c}$ denotes input embedding and $\beta$ is a trade-off parameter.
Intuition: The proofs are provided in Appendix A. The theoretical basis of MemVR draws support from the DPI (Cover et al., 1991) and the contraction properties of stochastic mappings in deep neural networks, as evidenced in numerous IB-related studies (Achille & Soatto, 2018). By enhancing MI and reducing uncertainty in hidden states, MemVR effectively alleviates hallucination while maintaining efficiency.